OpenStudy (anonymous):

Solve the system by elimination method. 0.5x=-2.0+y , -4/3x+y=-29/3

OpenStudy (anonymous):

do you know what the elimination method is?

OpenStudy (anonymous):

@marixoxo76

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

the elimination method is basically getting rid of one of the variables and solving

OpenStudy (anonymous):

the way we do this is to get both the x and y tems on the same side. after doing this what do you get?

OpenStudy (anonymous):

@marixoxo76

OpenStudy (anonymous):

i really don't understand with this problem..

OpenStudy (anonymous):

to get the variables to one side subtract a y from both sides of the equation like this 0.5x-y=-2.0 , -4/3x+y=-29/3

OpenStudy (anonymous):

now add these 2 equations together and you will see that the y variables cancel eachother out. then solve for x

OpenStudy (anonymous):

but how do you add if the second equation is a fraction?

OpenStudy (anonymous):

find a common denominator